### v1

# MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM

Abstract
Large language models (LLMs) have advanced code generation from single-function tasks to competitive-programming problems, but existing multi-agent solutions either rely on costly large-scale (>30B>30B) models or collapse when downsized to small open-source models. We present MapCoder-Lite, which upgrades a single 7B model into four role-specialised agents—retriever, planner, coder, and debugger—using only rank-32, role-specific LoRA adapters ($3\%$ extra parameters). Three lightweight techniques make this possible: (i) trajectory distillation from strong LLMs fixes format fragility in retrieval and debugging, (ii) supervisor-guided correction strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation. Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, and closes to within six points of a 32B baseline while cutting GPU memory and token-generation time by $4\times$. These results demonstrate that careful agent-wise fine-tuning unleashes high-quality multi-agent coding on a small language model.
MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM
Woongkyu Lee1   Junhee Cho2   Jungwook Choi1thanks: Corresponding author. 1Hanyang University  2Samsung SDS {lwghanyang, choij}@hanyang.ac.kr junhee.cho@samsung.com
1 Introduction
LLMs have revolutionized code synthesis, achieving near-perfect accuracy on function-level tasks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Research has now shifted to competitive programming, which demands efficient algorithms, robust implementation, and resilience against hidden test cases. Benchmarks like CodeElo Quan et al. (2025) highlight significant challenges that have spurred the development of massive models, such as OpenAI’s o3 OpenAI et al. (2025) and the 671 B-parameter DeepSeek-V3 DeepSeek-AI et al. (2025b), which dominate leaderboards despite their computational costs and proprietary nature.
To meet the cognitive breadth of competitive programming, recent research has shifted from monolithic prompting to multi-agent code-generation frameworks that allocate distinct roles to specialised language-model agents Islam et al. (2024). These systems separate problem solving into discrete stages—algorithm selection, planning, coding, and debugging—enabling focused expertise and structured information exchange between agents Huang et al. (2024); Hong et al. (2024); Islam et al. (2024, 2025). MapCoder Islam et al. (2024) (Fig. 1) exemplifies this approach with a pipeline where specialized agents handle different aspects of problem solving, narrowing the gap between open-source and proprietary solutions.
However, multi-agent systems present two significant challenges. First, they are computationally intensive when using large models. Our MapCoder implementation with Qwen-2.5-32B on the CodeContests benchmark required 3,000 forward passes, each involving 70 GB of parameter loading, resulting in 27.5 hours of runtime on a single A100 GPU—far exceeding the cost of single-prompt approaches. Second, downsizing the agents to 7B models leads to a sharp decline in reliability (see Fig. 2): (i) format adherence breaks down, with frequent XML schema violations that halt pipeline execution; and (ii) role-specific performance suffers, yielding brittle algorithms, incorrect code, and ineffective debugging. Achieving a balance between competitive performance and computational efficiency remains a key open problem.
Figure 1: Overview of the MapCoder system. Given a natural language problem, the retrieval agent fetches relevant algorithmic knowledge, followed by the planning agent generating a solution plan. The coding agent implements the plan, and the debugging agent iteratively refines the code based on test outcomes.
We address these challenges with MapCoder-Lite, the first multi-agent coding framework that drives a single 7B backbone–extended only by lightweight, role-specific adapters–to performance near that of 32B systems. MapCoder-Lite is built on three components:
  • Trajectory distillation from strong LLMs (Sec. 3.1): For retrieval and debugging roles—where small models often fail—we harvest high-quality traces from Qwen-32B and DeepSeek-V3, filtered through execution tests to ensure format correctness and semantic accuracy.
  • Supervisor-aided cross-agent refinement (Sec. 3.2): For planning and coding roles, we employ a supervisor model that identifies failures, provides targeted feedback, and regenerates problematic steps, creating high-quality training examples without human annotation.
  • Memory-efficient LoRA specialization (Sec. 3.3): All agents share a frozen Qwen-2.5-7B backbone with role-specific low-rank adapters (LoRA Hu et al. (2021)), adding <3% of the parameters of a 7B model while enabling independent fine-tuning.
We conducted a comprehensive evaluation on three representative competitive-programming suites—xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, and narrows the gap to a 32B MapCoder to within six points, all while cutting GPU memory and token generation time(time per output token) by $4\times$. These results demonstrate that careful agent-wise fine-tuning can unlock high-quality multi-agent code generation on small language models.
2 Background
2.1 LLMs for Competitive Programming
Large language models have already revolutionised function-level code generation: specialised variants such as Codex, Qwen-2.5-Coder, DeepSeek-Coder, and CodeLlama routinely achieve near-perfect scores on HumanEval, MBPP, and LiveCodeBench Chen et al. (2021); Hui et al. (2024); Guo et al. (2024); Rozière et al. (2024); Jain et al. (2024). But recent community attention has shifted toward the far more demanding arena of competitive programming, where a model must translate a prose problem statement into a complete, time- and memory-efficient program that solves all hidden test cases. Datasets such as APPS Hendrycks et al. (2021), CodeContests Li et al. (2022), and xCodeEval Khan et al. (2024) embody this shift: they assess full-program synthesis, algorithm selection, and constraint reasoning under the stringent standards of contest judging. Progress on these benchmarks now defines the “state of the art” for code-capable LLMs OpenAI et al. (2025); DeepSeek-AI et al. (2025a), but their difficulty also exposes the limits of single-agent prompting, motivating the multi-agent systems discussed next.
2.2 Multi-Agent Code Generation
As competitive programming problems grow exceptionally in linguistic length and algorithmic subtlety, it is not uncommon for a single LLM prompt to miss hidden edge cases or violate the constraints of the competition. Recent work, therefore, turns to multi-agent pipelines, in which a set of specialised LLM agents collaborate under an explicit workflow Huang et al. (2024); Pan et al. (2025). Among these systems, MapCoder Islam et al. (2024) stands out with its attractive design that performs well in APPS and CodeContests. Islam et al. (2024).
Fig. 1 illustrates MapCoder’s four-stage coding pipeline. A retrieval agent first queries a private corpus of textbook algorithms and returns a short, schema-constrained XML snippet that explains a candidate approach. Next, a planning agent expands this snippet into several step-wise execution plans, and a confidence score is generated for each plan using XML format. Plans are sorted by confidence, and the top one is passed to a coding agent that writes full source code in the required language. Finally, a debugging agent compiles, executes, and patches the program until it clears all unit tests; if it fails, the loop backtracks to the next plan.
Role specialisation makes the pipeline modular: each agent faces a simpler, well-scoped sub-task, so errors can be isolated and corrected step-by-step, yielding higher overall accuracy than a monolithic prompt can achieve. The price of this modularity, however, is that the underlying LLM must possess enough general capability to perform four very different roles—retrieving algorithms, reasoning about plans, writing code, and fixing bugs—without losing coherence across stages.
2.3 Task-Specific Fine-Tuning
MapCoder relies on large-scale (>30B>30B), general-purpose LLMs at every stage, trading speed and memory for accuracy. Because each agent tackles a well-defined sub-task—retrieving algorithms, planning steps, writing code, or fixing bugs—a natural question is whether a single 7 B backbone can match MapCoder if it is fine-tuned per role. Fine-tuning smaller models has proved effective in other multi-agent settings Zhao et al. (2025); Liang et al. (2025); Shen et al. (2024), and several studies have enhanced code-related abilities through supervised updates or parameter-efficient adapters Fan et al. (2025); Tsai et al. (2024); Yu et al. (2024); Jiang et al. (2024a). To our knowledge, however, no work has comprehensively studied how far role-aligned fine-tuning can push a small LLM inside a multi-agent pipeline for competitive programming.
Model Metric Cost
Number of problems 165
Total runtime 27.53 hours
Qwen32B Input tokens 5.08M tokens
Output tokens 1.60M tokens
Number of API calls 3,095 calls
Table 1: Resource cost of MapCoder using Qwen2.5-32B-Instruct on the CodeContests benchmark.
Figure 2: Representative failure cases in 7B-scale models. (a) Format following error in the retrieval agent: the model fails to generate responses in the required XML structure, leading to format mismatches and processing errors. (b) Incomplete reasoning in the planning agent: a key step is omitted from the solution plan, resulting in failure to fully cover possible cases.
2.4 Challenges
We identify key challenges in multi-agent code generation based on our implementation of MapCoder.
High Cost with Large Models. We evaluated MapCoder using the Qwen2.5-32B-Instruct model on the CodeContests benchmark, which comprises 165 competitive programming problems. As shown in Table 1, the system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls. This substantial resource usage stems from MapCoder’s multi-agent design involving four distinct agents and multiple iterations for planning and debugging. These results highlight the heavy runtime and memory burden of using large-scale language models (LMs) throughout the pipeline. We hypothesize that replacing each agent with a fine-tuned small LM can significantly reduce token-generation time and GPU memory usage, even with similar token and API call counts.
Failure Cases with Small Models. Fig. 2 illustrates typical failures when downsizing MapCoder’s model to a 7B scale.
  • Format Following Failures: Multi-agent systems often rely on structured outputs in predefined formats. For instance, in MapCoder, the retrieval and planning agents must return XML-formatted responses with tags like <root>, <algorithm> and <confidence>. In Fig. 2(a), the 7B model inserts an extraneous <description> tag and omits the closing </root> tag, violating the expected format. Such structural errors disrupt downstream processing, as subsequent agents or parsers depend on well-formed inputs. These issues are particularly critical in multi-agent pipelines, where format consistency ensures reliable coordination.
  • Low Role Performance: Small models often struggle with role-specific tasks due to limited capacity. As shown in Fig. 2(b), the 32B model’s planning agent enumerates all relevant cases, demonstrating robust and comprehensive reasoning. In contrast, the 7B model omits a key sub-case involving the combination of multiple ticket types. Incomplete planning like this can yield incorrect solutions, even if individual steps seem plausible. Similar shortcomings appear in other roles: the coding agent may generate incorrect logic or indexing, while the debugging agent frequently fails to identify and correct bugs, leading to repeated failures across iterations.
To mitigate these issues, we propose task-specific fine-tuning to improve both format compliance and role execution, thereby enhancing overall code generation accuracy.
3 Methodology
To replace MapCoder’s 32B backbone with a memory-efficient 7B model while maintaining accuracy, we introduce a role-aligned supervised fine-tuning (SFT) pipeline that tailors one shared 7B backbone to four agents—retrieval, planning, coding, and debugging—via lightweight LoRA adapters. These changes present three specific challenges:
  • Role-specific data scarcity: Public code-generation datasets provide ground-truth programs but lack datasets for retrieval queries, step-wise plans, and debug traces, hindering agent-wise supervision.
  • Cross-agent consistency: Naïvely fine-tuning each agent in isolation overlooks inter-agent dependencies; a fix applied to one step can still fail downstream.
  • Parameter budget inflation: Even if fine-tuning recovers agent-specific capability, storing four independent 7B checkpoints nullifies memory savings, approaching the original 32B footprint.
We address (i) by distilling high-quality trajectories from a strong LLM for retrieval & debugging, tackle (ii) with a supervisor model that diagnoses end-to-end failures and regenerates only the faulty segment for planning & coding, and solve (iii) by sharing a single 7B backbone augmented with rank-32 LoRAs—adding $3\%$ parameters while preserving MapCoder’s modularity.
Figure 3: Illustration of trajectory construction for retrieval and debugging datasets. (a) When a strong LLM generates a complete solution that passes unit tests, the retrieval agent’s input-output pair is extracted as a training example. (b) To collect debugging data, a 7B model is used for planning and coding, and the strong LLM is used for debugging when the initial code fails. If the revised output passes, the debugging trajectory is added to the dataset.
3.1 Strong LLM for Retrieval and Debugging
Observations. Two weaknesses repeatedly surface when the 32B backbone is replaced with a 7B model. First, the retrieval agent often violates the mandated XML schema or returns partial algorithm descriptions, so downstream agents cannot even parse its output. Second, the debugging agent fails to patch seemingly simple boundary-case bugs, leaving many programs one step short of success. Experiments with stronger LLMs such as Qwen2.5-32B-Instruct and DeepSeek-V3 show that these larger LLMs seldom make such mistakes, indicating that their trajectories can serve as high-quality supervision for smaller LLMs.
Challenges. Effective role-specific fine-tuning is hampered by two data gaps. (i) Public code-generation benchmarks provide only problem–solution pairs and contain no retrieval queries, step-wise plans, or debug traces; thus, no off-the-shelf corpus targets the intermediate process of the multi-agent coding pipeline. (ii) Even when synthetic trajectories are generated, many are noisy and fail the unit tests, so naïvely training on them would propagate errors rather than correct them.
Methods. Fig. 3 shows the proposed trajectory construction method. Our data pipeline begins by asking strong LLMs—Qwen2.5-32B-Instruct or DeepSeek-V3—to solve each coding task while explicitly printing the intermediate artefacts (i.e. trajectories) obtained from each MapCoder role. In one shot, the model emits an algorithm description for the retrieval agent, a step-wise plan for the planning agent, source code for the coding agent, and a patched version for the debugging agent. Because the same LLM generates the entire trajectory, the four outputs are mutually consistent and naturally aligned with their respective prompts.
To convert these raw trajectories into reliable supervision, we exploit the fact that MapCoder’s final objective—passing all the unit tests—is objectively verifiable. After executing the LLM-produced code, we keep the complete trajectory only if the program passes. This pass‐based filtering eliminates noisy or partially correct traces without any manual inspection: a failure at test time almost always implies that at least one upstream artefact (faulty retrieval cue, flawed plan, or ineffective patch) is wrong, so discarding the whole sequence prevents error propagation during fine-tuning. The surviving set—about 2,200 retrieval examples and 4,200 debugging examples from APPS, xCodeEval, and CodeContests—forms a high-quality, role-aligned corpus that directly trains rank-32 LoRA on the shared 7B backbone. Empirically, models tuned on this filtered dataset regain XML-schema fidelity and bug-repair accuracy comparable to the original 32 B system while retaining the memory footprint advantages of a 7B model (see Table 2 for resulting accuracy enhancement).
Figure 4: Supervisor-aided data collection pipeline. When the final output fails, the supervisor inspects the full trajectory (including algorithm, plan, code, and test result), identifies the responsible agent, and provides targeted feedback to revise its output. If the revised result passes, the updated trajectory is added to the fine-tuning dataset.
Retrieval Planning / Coding xCodeEval
Dataset
Pass
Filtering
Dataset
Pass
Filtering
Accuracy (%)
11.32
Strong 16.04
Strong 16.98
Strong Strong 18.87
Strong Supervisor 22.64
Table 2: Ablation study on data source and pass-based filtering for training retrieval, planning, and coding agents on xCodeEval. “Strong” refers to data generated by a large LM (e.g., Qwen2.5-32B, DeepSeek-V3), and “Supervisor” denotes data refined by a supervisor model after failure analysis.
3.2 Supervisor-Guided Planning and Coding
Observations. Even after the role-specific fine-tuning on strong LLM trajectories, the 7B planning agent still emits fragmentary execution flows, omitting edge cases, misordering steps, or drifting from the prescribed outline. Also, the coding agent continues to generate programs that compile yet collapse under the full unit-test suite. These local errors propagate: an incomplete plan starves the coder of crucial hints, and a superficially correct program deceives the debugger until final evaluation. Table 2 confirms that this trajectory-only adaptation closes part of the gap but suffers from a noticeable accuracy loss. In contrast, Qwen2.5-32B-Instruct and DeepSeek-V3 produce structurally sound plans and largely correct first-pass implementations, indicating that stronger models possess global cues that the smaller model has not yet internalised.
Challenges. Despite the availability of agent-wise strong-model trajectories, two obstacles remain for high-quality fine-tuning. First, a locally tuned agent lacks global awareness: improving its own output does not guarantee end-to-end success unless that output is judged in context. Second, the capacity gap hinders direct imitation: a 7B model exposed to large-scale models’ plans and coding results may underfit or copy surface patterns without learning the deeper reasoning (phenomenon called capacity gap Bansal et al. (2024); Xu et al. (2025)).
Methods. We propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model. As shown in Fig. 4, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory. If the resulting program fails its unit tests, the entire trajectory is forwarded to a high-capacity supervisor LLM (DeepSeek-V3). The supervisor analyses the trace, pinpoints the agent primarily responsible for the failure, and issues concise, role-specific feedback, and only that agent regenerates its output. The revised trajectory is re-tested; once all tests pass, the final plan–code pair is added to the fine-tuning corpus.
Supervisor-guided refinement is applied selectively to trajectories where the strong LLM succeeds but the 7B model fails, minimizing generation overhead. The supervisor operates only during data generation, keeping inference lightweight. We store just the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime. Crucially, every example in this corpus is produced and execution-validated by the 7B model, eliminating concerns about a capacity gap. Iterating over thousands of problems yields a dataset that aligns planners and coders with end-to-end success. Table 2 confirms that this strategy delivers a substantial accuracy boost.
FT Method
Trainable
Parameters
Required Memory
for Training (GB)
Accuracy
FFT 7615.62M 45.69 18.87
LoRA 20.19M 15.35 22.64
Table 3: The number of trainable parameters, required memory for training, and xCodeEval accuracy of full fine-tuning (FFT) and LoRA methods with MapCoder-Lite 7B. Accuracy is measured only up to the coding stage, without debugging.
3.3 Multi-Agent LoRA Fine-Tuning
To enable role-specific specialization in our multi-agent framework, we adopt an agent-wise LoRA approach. All four agents—retrieval, planning, coding, and debugging—share a frozen 7B base model, each paired with its own lightweight LoRA. These adapters are fine-tuned independently using pass-based-filtered and supervisor-guided trajectory distillation. This setup significantly reduces GPU memory usage during fine-tuning and supports modular training and deployment, allowing each agent to be updated without duplicating the full model. As shown in Table 3, LoRA significantly reduces the number of trainable parameters and memory footprint, while achieving even higher accuracy than full fine-tuning. We attribute this accuracy gain in part to LoRA’s implicit regularization, which steers the model toward role-specific behaviors without overwriting core pretrained knowledge—a finding consistent with prior work Hu et al. (2021); Houlsby et al. (2019).
4 Experimental Settings
Method xCodeEval (106) APPS (150) CodeContests (165)
[…]
Direct 18.87 20 4.00 6 1.21 2
CoT 5.66 6 4.67 7 3.64 6
Self-planning 9.43 10 1.33 2 3.03 5
Analogical 12.26 13 3.33 5 1.82 3
MapCoder 13.21 14 29 6.00 9 30 6.06 10 59
MapCoder-Lite 28.30 29 0 8.00 12 0 13.33 22 0
Table 4: Performance comparison of different methods using Qwen7B on xCodeEval, APPS, and CodeContests with Pass@1 accuracy (%). The numbers in parentheses indicate the number of problems in each benchmark.
Models. All agent-wise fine-tuning experiments use Qwen2.5-7B-Instruct as the base model. For comparison, we include larger models—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, serving as upper bounds to assess the effectiveness of our lightweight strategy.
Benchmarks. We evaluate models on three competitive programming benchmarks: xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022). For xCodeEval, we use the 106 problems from the compact program synthesis split. For CodeContests, we evaluate all 165 official test problems. We randomly select 150 test-set problems for APPS, following MapCoder Islam et al. (2024). Each benchmark includes natural language problem descriptions with sample I/O and is evaluated using hidden unit tests. We also optionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
Training Configuration. Each agent is fine-tuned using LoRA adapters with rank 32 applied to the linear modules for query, key, value, and output projection. Training is implemented with Hugging Face Transformers and PEFT libraries HuggingFace , using a learning rate of 2e-5, gradient accumulation of 16, and 3 training epochs. Hyperparameters were tuned individually per agent to ensure stable and consistent performance. All training experiments were conducted on a single NVIDIA A100 80GB GPU.
Evaluation Setting. We report pass@1 accuracy as the primary metric, reflecting whether the model’s top-1 output passes all test cases. All outputs are generated using greedy decoding for reproducibility and consistency with baseline evaluation.
5 Experimental Results
5.1 Effectiveness of Agent-wise Fine-Tuning
Table 4 compares the performance of a small LM (Qwen2.5-7B) across various single-agent prompt optimization methods—direct prompting, chain-of-thought (CoT) Wei et al. (2023), self-planning Jiang et al. (2024b), and analogical prompting Yasunaga et al. (2024)—as well as multi-agent approaches: MapCoder and our proposed MapCoder-Lite. Without fine-tuning, MapCoder shows only marginal gains over single-agent methods and even underperforms Direct prompting on xCodeEval (13.21% vs. 18.87%), despite using four specialized agents. This highlights the limitations of small LMs in multi-agent setups, where poor coordination and frequent format violations (as indicated by numerous format failures) prevent reliable end-to-end performance.
In contrast, MapCoder-Lite, with agent-wise fine-tuning, achieves substantial improvements across all benchmarks. On xCodeEval, accuracy more than doubles to 28.30%, outperforming all prompting baselines. Similar gains are seen on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%). Notably, fine-tuning eliminates all format failures, indicating that even small LMs can produce structurally consistent outputs when role-specialized. These results emphasize the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks.
5.2 Comparison with Larger Models
Table 5 compares MapCoder’s performance across backbone sizes from 7B to 32B (Qwen[7–32]B), alongside MapCoder-Lite—our fine-tuned 7B variant (Qwen7B-FT). Qwen7B-FT delivers a substantial accuracy boost over the untuned Qwen7B, matching the performance of Qwen14B and coming within six points of Qwen32B. This is particularly notable given that Qwen7B-FT requires only one-quarter of the GPU memory of Qwen32B, making it viable for deployment on memory-constrained edge devices Karumbunathan (2022). Moreover, the reduced model size yields a proportional speedup in token generation—LLM decoding being memory-bound—achieving a 4× reduction in time-per-output-token (TPOT), as measured using vLLM Kwon et al. (2023). These results highlight that targeted fine-tuning enables small models to achieve competitive performance with dramatically lower cost.
Model xCodeEval APPS CodeContests Memory TPOT
Qwen7B 13.21 6.00 6.06 15.26 12.29
Qwen7B-FT 28.30 8.00 13.33 15.64 12.29
Qwen14B 28.30 10.00 15.76 29.58 23.36
Qwen32B 33.02 13.33 18.18 65.56 45.06
Table 5: Pass@1 accuracy (%) of MapCoder using Qwen models of different scales on xCodeEval, APPS, and CodeContests. Memory usage (GB) and Time Per Output Tokens (ms).
Agent xCodeEval Pass Pass Format
[…]
- - - - 13.21 11 3 29
- - - 18.87 17 3 12
- - 26.42 18 10 0
- 24.53 24 2 0
- 27.36 17 12 0
28.30 24 6 0
Table 6: Ablation study of per-agent fine-tuning contributions in MapCoder on xCodeEval. R, P, C, and D denote Retrieval, Planning, Coding, and Debugging agents, respectively.
Method HumanEval (164 problems) MBPP (397 problems)
[…]
MapCoder 70.73 105 11 14 67.51 244 24 68
MapCoder-Lite 82.93 120 16 1 84.63 305 31 0
Table 7: Evaluation of MapCoder and MapCoder-Lite on HumanEval and MBPP benchmarks.
5.3 Ablation Study
Contribution of Individual Agent Fine-Tuning. Table 6 shows an ablation study evaluating the impact of fine-tuning each agent in MapCoder on xCodeEval. Fine-tuning the retrieval agent alone improves accuracy from 13.21% to 18.87% and reduces format failures from 29 to 12, indicating better structural consistency. Adding the planning agent further boosts accuracy to 26.42% and eliminates format errors entirely, increasing both non-debug passes (17 → 18) and debug-assisted completions (3 → 10). Fine-tuning the coding agent (without debugging) raises non-debug passes to 24 but reduces debug-assisted passes to 2, since MapCoder only invokes the debugger when initial code fails sample I/O tests. Stronger code may bypass debugging entirely, leaving some errors uncaught. Conversely, fine-tuning the debugging agent (without coding) increases accuracy to 27.36%, with 12 debug-assisted completions—showing that an effective debugger can recover from weaker initial code. When all four agents are fine-tuned, overall accuracy peaks at 28.30%, with a strong balance between direct (24) and debug-assisted (6) completions. These results demonstrate that each agent addresses distinct failure modes and that coordinated fine-tuning is essential for maximizing end-to-end performance.
Evaluation on Simpler Coding Tasks. To evaluate generalization beyond competition-level programming, we compare MapCoder and MapCoder-Lite on two function-level benchmarks: HumanEval and MBPP. As shown in Table 7, MapCoder-Lite outperforms the baseline by significantly reducing format-related failures and achieving higher end-to-end accuracy. It improves both coding (increased passes without debugging) and debugging (even greater gains in passes with debugging), indicating enhanced reliability. These results suggest that MapCoder-Lite not only preserves generalization capability but also improves robustness on structurally simpler tasks.
Multi-Agent Tuning with Coder Model. We evaluate whether multi-agent code generation can be enhanced using a code-specialized language model (a.k.a. Coder model), such as Qwen2.5-Coder-7B-Instruct Hui et al. (2024). As shown in Table 8, applying the multi-agent framework (MapCoder) and our fine-tuning strategy (MapCoder-Lite) to the Coder model does improve its performance, demonstrating that even domain-specialized models benefit from structured, role-specific tuning. However, despite these gains, the Coder-based pipeline still lags significantly behind our baseline model, Qwen2.5-7B-Instruct, both before and after fine-tuning. We attribute this gap to the Coder model’s limited general reasoning ability, which is critical for handling the diverse roles in multi-agent workflows, including retrieval, planning, and debugging, in addition to coding. In contrast, the general-purpose Qwen2.5-7B-Instruct model more effectively supports multi-agent coordination, and with our memory-efficient fine-tuning (MapCoder-Lite), achieves even higher end-to-end accuracy.
Model Method Accuracy (%)
Qwen2.5-Coder-7B-Instruct MapCoder 5.66
MapCoder-Lite 15.09
Qwen2.5-7B-Instruct MapCoder 11.32
MapCoder-Lite 22.64
Table 8: Comparison of MapCoder and MapCoder Lite using Qwen2.5-Coder-7B-Instruct. Accuracy is measured only up to the coding stage, without debugging.
5.4 Case Study
To qualitatively assess the impact of agent-wise fine-tuning, we examine representative before-and-after examples for each agent. As illustrated in Appendix B, fine-tuning leads to notable improvements in algorithm retrieval, plan coverage, code correctness, and bug resolution. These examples highlight how each agent benefits from role-specific supervision, supporting the effectiveness of our fine-tuning strategy.
6 Conclusion
We introduced MapCoder-Lite, a multi-agent coding framework built on a 7B model with lightweight LoRA adapters. Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance with lower cost. Our results show that small models, when carefully fine-tuned, can match the reliability and accuracy of larger systems.
Limitations
Our framework relies on distilled trajectories from strong models like DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance. Rather than indicating a ceiling, this gap highlights opportunities to further enhance small models—potentially through architectural extensions, better training objectives, or reinforcement-based learning from distillation feedback.
In addition, the multi-agent structure introduces a wide design space for fine-tuning, as each agent has a distinct role and optimization goal. While we adopted a uniform tuning recipe across agents, this modularity also enables targeted specialization and co-training strategies in future work, suggesting that more adaptive tuning could unlock even greater performance.
Appendix A Details of Supervisor
Fig. 5 and Fig. 6 illustrate the operation of the supervisor used in our data curation pipeline. The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, plan, code), and the result of test case execution.
As shown in Fig. 5, the input prompt explicitly instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent. The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that can be used to re-generate only the faulty component.
Fig. 6 shows a typical response from the supervisor. It correctly attributes the failure to the planning agent, explains why the plan is insufficient, and suggests how it should be modified. This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without requiring human annotation.
Appendix B Improvements After Fine-Tuning
Fig. 789, and 10 illustrate representative improvements observed in the 7B model for the retrieval, planning, coding, and debugging agents, respectively.
In the retrieval example (Fig. 7), the pre-fine-tuned model incorrectly identifies the core algorithm as “Sliding Window” and includes unsupported tags like <description> within the <algorithm> block, failing to conform to the expected XML schema. After fine-tuning, the agent accurately identifies the algorithm (“Counting and Matching Pairs”) and outputs a well-structured XML response with all required tags properly closed.
In the planning example (Fig. 8), the original plan fails to capture key logical conditions (e.g., that the input must be both even and greater than 2). After fine-tuning, the agent successfully generates a plan that handles these conditions explicitly and correctly guides the coding agent.
In the coding example (Fig. 9), the unfine-tuned model incorrectly processes multiple input values on a single line, violating the problem’s input specification. Fine-tuning enables the agent to read inputs line by line, aligning its behavior with the expected input format and improving functional correctness.
Finally, the debugging example (Fig. 10) demonstrates how the unfine-tuned model fails to fix a bug caused by improper input parsing. The fine-tuned debugging agent correctly diagnoses the issue and proposes a revised plan and code that successfully passes all test cases.
These examples highlight the role-specific benefits of fine-tuning, improving not only task-level correctness but also structural and logical alignment across the entire pipeline.
Figure 5: Prompt for the Supervisor.
Figure 6: Example response from the supervisor: identifies the flawed part of the plan and generates feedback.
Figure 7: Improvement in algorithm tutorial and XML formatting by the retrieval agent after fine-tuning. The pre-trained model fails to identify the correct algorithm and generates an ill-formed XML response with an unclosed <root> tag. After fine-tuning, the retrieval agent provides a valid explanation based on matching command pairs and generates a well-structured XML block.
Figure 8: Improvement in conditional logic by the planning agent after fine-tuning. The initial code fails to account for the edge case (w > 2) and only checks for evenness. After fine-tuning, the planning agent correctly generates the full condition required by the problem.
Figure 9: Example of a coding error and its resolution after fine-tuning. The unfine-tuned agent incorrectly reads all input values from a single line, which violates the problem specification. After fine-tuning, the coding agent correctly processes line-separated inputs and computes the result accordingly.
Figure 10: Example of a debugging failure and recovery in the MapCoder pipeline. The initial plan and code fail to address the problem due to incorrect input parsing. After the debugging agent identifies the issue, the revised plan introduces line-by-line input reading, which successfully resolves the error.